# DAViD: Modeling Dynamic Affordance of 3D Objects using Pre-trained Video Diffusion Models

### v2

@@ 1 Introduction
Humans effortlessly understand how to effectively use objects to achieve their goals within a given environment. This knowledge, known as affordance, originally introduced by Gibson [21], encapsulates a comprehensive understanding of the inherent ways humans interact with objects. Modeling these patterns is crucial for AI systems to assist or mimic human behavior, leading to extensive research in AI fields [93, 85, 79, 92, 57, 72, 40, 1, 63, 61, 64, 78, 65, 15, 10, 121, 9, 113, 123, 97, 53, 62, 71, 122, 12, 75, 36, 55, 11]. However, existing studies on affordance learning primarily focus on static human-object interaction (HOI) patterns. For example, some approaches address affordances by inferring contact (or interaction) region between humans and objects in 2D images [49, 2], 3D humans [90, 110], and 3D objects [109, 110]. Others [26, 37] represent affordances as spatial distributions aggregated from static 3D HOI samples. While recent methods generate human motions including HOI [50, 107, 70, 51, 108, 105, 45] and Human-Scene Interaction (HSI) [100, 112, 105], they often focus on limited actions such as standing, sitting, or picking, without focusing on rich, dynamic nature of object-specific affordances in human motion.
@@
In this paper, we present a novel framework for learning Dynamic Affordance across diverse and unbounded target object categories, without the need for laborious 3D HOI data capture. Our approach is motivated by the observation that human-object interactions involve not only static properties (e.g., contact regions or grasping patterns) but also object-specific characteristic movements over time. For instance, as shown in Fig. 2, humans naturally push a shopping cart and pull a suitcase while walking forward, yet rarely lift their handles vertically. Existing studies, which mainly focus on static affordance features, fail to capture these dynamic interaction patterns. To address this, we present DAViD, a generative motion model for modeling Dynamic Affordance that learns the dynamic and object-specific 4D HOI patterns. To handle the vast diversity of dynamic HOI patterns, our method considers each object-specific 4D HOI pattern as a distinct motion concept, adapting a pre-trained motion diffusion model (MDM) [88] via Low-Rank Adaptation (LoRA) [31]. Once human motions are generated by our MDM with LoRA, we further synthesize object motions conditioned by human motion using a score-based diffusion model, where we introduce a novel sampling technique, named Temporal Guidance Sampling (TGS), for enhancing motion consistency and fine details. Notably, our approach goes beyond modeling dynamic affordance with limited data samples, enables to naturally blend the newly learned HOI motion concepts with motions already known from the pre-trained model in a zero-shot manner, even for multiple HOI concepts.
@@
To train DAViD, we also present a pipeline to synthetically produce realistic 4D HOI samples for a target object category, enabling effective model training without requiring expensive real-world 3D data collection. Most existing dynamic HOI learning approaches [98, 33, 116] rely on the 3D datasets captured in lab-controlled environments [3, 32, 28, 119, 86, 18, 38]. This constraint significantly limits the opportunity to learn diverse and distinctive dynamic affordance patterns across a broader range of object categories. Our key insight is that pre-trained video generation models inherently capture dynamic patterns of object usage. As shown in Fig. 2, these models already know plausible dynamic interactions (e.g., pushing a shopping cart or pulling a suitcase). Since it presents non-trivial challenges to leverage the synthetic 2D videos for learning 4D dynamic affordance, we propose a novel pipeline that synthesizes plausible 4D HOI samples for a given 3D object mesh. We achieve this by generating 2D videos aligned to the 3D object, and aligning the uplifted 4D human motion sequences with the object model, guided by synthetic video cues.
@@
We validate the efficacy of our approach by learning Dynamic Affordance on 30 3D object categories, gathered from multiple sources [3, 32, 8, 106, 60, 37, 83]. We compare our DAViD with other HOI motion synthesis approaches [50, 51, 88] against public dataset [50], demonstrating that our model outperforms competitors in generating HOI motion with plausible contact. We further demonstrate the advantage of our pipeline, by generating results for a various previously unexplored target objects, and also validating its advantage in merging multiple HOI motion concepts with known humn motion from pre-trained MDM [88].
@@
In summary, our main contributions are as follows: (1) We present a generative motion model, DAViD, to learn Dynamic Affordance; (2) We present a pipeline to synthetically generate diverse 4D HOI samples leveraging video diffusion models; (3) We address generating various novel HOI motions by fusing multiple, separately trained LoRAs with known human motions from a pre-trained model. The results and our source code will be released.
@@ 2 Related Work
Reasoning Visual Affordance. Affordance is a concept defined by Gibson [21] as the set of possible interactions that an agent can perform within an environment. As the concept narrows to the primary agent (or humans) and the specific interaction target (or objects), the area of learning affordance mainly focuses on HOI. Specifically, followed by the intuition that the visual geometry of an object is closely related to its functionality, many studies leverage visual cues to reason about object’s affordance. The studies are subdivided into various areas, including affordance classification [93, 85, 79, 92, 57, 72, 40], affordance detection [1, 63, 61, 64, 78, 65, 15, 10], and affordance segmentation [75, 36, 55, 11]. Such traditional studies infer affordance for the given context (mostly provided as images), as a form of action labels [5, 30, 47], and contact (or interaction) regions in 2D [49, 2] or 3D [90, 109]. Later studies model static HOI patterns (e.g., contact) during interactions [13, 26, 37], going beyond the understanding of individual HOI situations. Recent studies synthesize human motion including HOI [107, 70, 100, 112, 51, 108, 105], showing potential to learn dynamic HOI patterns, but they focus on a text-driven approach, generating quite consistent human motions regardless of object categories (e.g., dragging a chair, holding a backpack). In contrast, our method focuses on learning rich, dynamic object-specific HOI patterns, observable in real-world HOIs.
Learning from Synthesized Data. Generative models such as GANs [22] and diffusion models are used as a strategy to address the problem of insufficient real-world data in many fields. Starting from augmentation [89, 29, 104, 66], synthetic datasets are widely used in vision area including segmentation [58, 91, 118], classification [7, 56, 87], shape reconstruction [68], and neural rendering [27, 117]. While data synthesis is a powerful technique that can be applied in many fields, the important thing to consider when learning from generated data is how reliable the generated data is compared to real-world data. Recently, few studies in the HOI field [26, 37] leverage pre-trained 2D diffusion model to generate 2D HOI images, demonstrating the reliability of synthetic data to learn affordance knowledge. While following the approaches, we focus on learning dynamic patterns during HOI rather than learning static patterns.
@@ 3 Method
Our approach aims to learn the Dynamic Affordance of a target object from synthetically generated 2D videos, which are produced by a video diffusion model [41]. In the first stage, we generate diverse 4D HOI samples interacting with a given 3D object mesh by leveraging pre-trained video models (Sec. 3.1). Then, our method models the collected 4D HOI samples into our DAViD, which consists of (1) a motion diffusion model along with LoRA, and (2) an object pose diffusion model. The first module synthesizes feasible human motions for the target object, and the second module synthesizes object motions paired with previously generated human motion, using our novel sampling technique, TGS (Sec 3.2). An overview of our method is shown in Fig. 3.
@@
2D HOI Video Synthesis. Given an input 3D object mesh, we produce multiple 2D HOI videos capturing human-object interactions with the object. We first render the object from cameras with multiple viewpoints. The intrinsic parameter is fixed to match the same perspective camera model and parameters of the world-grounded human mesh recovery method (GVHMR [81]) used in our later 3D lifting process, ensuring a consistent camera setup. For the extrinsic parameters, we consider diverse viewpoints based on the object’s type, either ground-placed object or portable. For relatively stationary object (e.g., cart), we consider horizontal viewpoint variations centered on the object, while for the portable objects (e.g., trumpet) we sample viewpoints within a specified range from all view angles. The object type is automatically inferred via a vision-language model [67], by querying with the rendered image with a prompt: “Is the object plausibly on the ground while being used?”.
Once the object is rendered from each view, we synthesize a realistic human naturally interacting with it using an off-the-shelf pre-trained 2D image diffusion model [44]. Unlike the previous similar approaches [42, 111, 37] that use inpainting by providing a mask region to a model, we use ControlNet [115] to synthesize the entire image, including the human and the plausible background, by using edge cues of rendered object [6] as the condition. Our primary objective is to synthesize both plausible human-object interaction for the given object viewpoint and a realistic background, which is important for estimating dynamic cues and camera poses in a global coordinate. After synthesizing HOI images from several views, we refine the results by filtering out low-quality or undesirable images. We find it is beneficial to use small amount of human effort for this, taking approximately three minutes per category.
@@
Figure 3: Overview. Our method consists of two parts: (1) Generating 4D HOI Samples and (2) Learning Dynamic Affordance from Generated 4D HOI Samples. First, we create 2D HOI videos, and generate 4D HOI samples with our uplifting pipeline. Then, the generated 4D HOI samples are used to train DAViD, learning the patterns of human motion and human pose conditioned object pose.
Uplifting for 4D HOI Sample Generations. For each 2D HOI video $\mathcal{V}$ 222we drop the video index $d$ for simplicity., we uplift 3D human motions and 3D object movements to generate 4D HOI samples. Our key strategy is to uplift each component (human, object) by finding each of the relative camera poses over time, and co-locate them in the common camera coordinate shared by both components.
For the object side, we use Perspective-n-Points (PnP) [19, 48] algorithm to find the relative camera pose with respect to the given 3D object mesh. Specifically, we aim to compute the extrinsic camera poses $\{\bm{R}_{o}^{f},\bm{t}_{o}^{f}\}_{f=1}^{F}$, where the one for the first frame $f=1$ is already known. We first identify visible object mesh vertices $\{\mathbf{V}_{o}\}_{\text{vis}}$ via raycasting [74] for the first frame $f=1$, where $\mathbf{V}_{o}\in\mathbb{R}^{3}$ is an object vertex. The 2D projections of the visible vertices at the first frame is denoted as $\{\bm{v}_{o}^{f=1}\}_{\text{vis}}$, where $\bm{v}_{o}^{f=1}\in\mathbb{R}^{2}$. By applying an off-the-shelf 2D video tracker [35, 34], we obtain the corresponding 2D points $\bm{v}_{o}^{f}$ in the remaining frames $f>1$.
As we have the correspondences between 3D object vertices and matched 2D points for every frame, we apply PnP [19, 48] to compute the camera poses at frame $f$ with respect to the object mesh in its canonical space:
@@
For the human side, we leverage world-grounded human motion recovery models, GVHMR [81], denoted as $\mathbf{F}_{\text{human}}$ to uplift human motions into 3D from the synthesized 2D HOI videos. We use GVHMR [81], except in the cases where human translation occurs without walking motions (e.g., riding a motorcycle). In such cases, we use TRAM [99], as GVHMR [81] typically fails to predict plausible camera motion under these conditions. Therefore,
@@
[…]
@@
where $\bm{V}_{h}^{f}\in\mathbb{R}^{3}$ is a human 3D vertex in its global coordinate after applying all SMPL parameters, and $\bm{\hat{V}}_{o}^{f}$ and $\bm{\hat{V}}_{h}^{f}$ are the human and object 3D vertices in the shared camera coordinate. To this end, we place both the 3D human and the 3D object in the global space of the human side (defined by GVHMR) to obtain their global 4D movements.
@@
where $\lambda$s denote the weighting factors. The $\mathcal{L}_{h}$ and $\mathcal{L}_{o}$ are the depth consistency loss for human and object, defined as:
@@
where $\bm{\hat{V}}_{h,i}$ and $\bm{\hat{V}}_{o,i}$ are the $i$-th human and object vertices in the camera coordinate at the first frame respectively, and $\bm{P}_{h,\text{3D}}$, $\bm{P}_{o,\text{3D}}$ are the corresponding 3D points by back-projecting 2D metric depths (computed by metric depth model [4]) into the camera coordinate. Note that the index $i,j$ are from the set of visible vertex indices of human and object at the first frame.
@@ 3.2 Learning Dynamic Affordance
Human Motion Synthesis with LoRA for MDM. To capture human motion patterns interacting with a specific object, we insert a LoRA [31] layer into a pre-trained MDM [88] model and train it for each object category. By leveraging LoRA [31], our model can learn previously unseen human motion patterns with a limited amount of training samples, keeping the original motion synthesis quality of the pre-trained model. Interestingly, we find that merging LoRA [31] layers trained accross different object categories enables our model to generate human motion interacting with multiple objects in a zero-shot manner (see Fig. 5 (c)).
@@ 4.2 Baselines and Metric
Even our goal is to learn dynamic HOI patterns for various objects, which is slightly different from typical text-driven HOI motion synthesis [14, 70, 108, 112, 107, 103], we conduct quantitative comparisons on existing datasets for evaluation. Specifically, we compare DAViD against OMOMO [50] and CHOIS [51] on FullBodyManipulation Dataset [50], which provide available code and evaluate the quality of HOI, unlike other stuides [14, 70, 103] which focus on evaluating human quality solely. However, OMOMO [50] and CHOIS [51] generate human motion conditioned on object motion (or waypoints), making a direct comparison with ours challenging. To address this, we compare the HOI motion generated by baselines using object motion (or waypoints) condition with the HOI motion generated by DAViD using human motion condition. Each condition is extracted from the test dataset of FullBodyManipulation [50], enabling a comparison with the ground truth. As using different condition, we compare only the quality of HOI for the fairness, specifically the precision, recall, and F1 score of the contact, following baselines [50, 51]. As DAViD performs modeling of object poses during interaction, we evaluate by cropping each test dataset sequence from the start of the interaction (contact start) to the end (contact end). Additionally, to assess the human motion quality, we report the Foot Sliding (FS) and foot height (in centimeters) of the human motion generated from DAViD and the baselines, both trained on FullBodyManipulation Dataset [50].

### v3

@@ 1 Introduction
Humans effortlessly understand how to effectively use objects to achieve their goals within a given environment. This knowledge, known as affordance originally introduced by Gibson [23], encapsulates a comprehensive understanding of the inherent ways humans interact with objects. Modeling these patterns is crucial for AI systems to assist or mimic human behavior, leading to extensive research in AI fields [41, 89, 83, 96, 60, 76, 43, 1, 66, 64, 67, 82, 68, 17, 12, 126, 11, 118, 128, 101, 56, 65, 75, 127, 14, 79, 38, 58, 13]. However, existing studies on affordance learning primarily focus on static human-object interaction (HOI) patterns. For example, some approaches address affordances by inferring contact (or interaction) region between humans and objects in 2D images [52, 2], 3D humans [94, 115], and 3D objects [114, 115]. Others [28, 39] represent affordances as spatial distributions aggregated from static 3D HOI samples. While recent methods generate human motions including HOI [53, 111, 74, 54, 112, 108, 48] and Human-Scene Interaction (HSI) [104, 117, 108], they often focus on limited actions such as standing, sitting, or picking, without focusing on rich, dynamic nature of object-specific affordances in human motion.
@@
In this paper, we present a novel framework for learning Dynamic Affordance across diverse and unbounded target object categories, without the need for laborious 3D HOI data capture. Our approach is motivated by the observation that human-object interactions involve not only static properties (e.g., contact regions or grasping patterns) but also object-specific characteristic movements over time. For instance, as shown in Fig. 2, humans naturally push a shopping cart and pull a suitcase while walking forward, yet rarely lift their handles vertically. Existing studies, which mainly focus on static affordance features, fail to capture these dynamic interaction patterns. To address this, we present DAViD, a generative motion model for modeling Dynamic Affordance that learns the dynamic and object-specific 4D HOI patterns. To handle the vast diversity of dynamic HOI patterns, our method considers each object-specific 4D HOI pattern as a distinct motion concept, adapting a pre-trained motion diffusion model (MDM) [92] via Low-Rank Adaptation (LoRA) [33]. Once human motions are generated by our MDM with LoRA, we further synthesize object motions conditioned by human motion using a score-based diffusion model, where we introduce a novel sampling technique, named Temporal Guidance Sampling (TGS), for enhancing motion consistency and fine details. Notably, our approach not only models dynamic affordances from limited data but also seamlessly blends newly learned HOI motion concepts with those already learnt in the pre-trained model, enabling zero-shot generalization, even across multiple HOI concepts.
@@
To train DAViD, we also present a pipeline to synthetically produce realistic 4D HOI samples for a target object category, enabling effective model training without requiring expensive real-world 3D data collection. Most existing dynamic HOI learning [102, 35, 121] or mimicking approaches [113] rely on the 3D datasets captured in lab-controlled environments [4, 34, 30, 124, 90, 20, 40]. This significantly limits the opportunity to learn diverse and distinctive dynamic affordance patterns across a broader range of object categories. Our key insight is that pre-trained video generation models inherently capture dynamic patterns of object usage. As shown in Fig. 2, these models already know plausible dynamic interactions (e.g., pushing a shopping cart or pulling a suitcase). Since it presents non-trivial challenges to leverage the synthetic 2D videos for learning 4D dynamic affordance, we propose a novel pipeline that synthesizes plausible 4D HOI samples for a given 3D object mesh. We achieve this by generating 2D videos aligned to the 3D object, and aligning the uplifted 4D human motion sequences with the object model, guided by synthetic video cues.
@@
We validate the efficacy of our approach by learning Dynamic Affordance on 30 3D object categories, gathered from multiple sources [4, 34, 10, 109, 63, 39, 87]. We compare our DAViD with other HOI motion synthesis approaches [53, 54, 92] against public dataset [53], demonstrating that our model outperforms competitors in generating HOI motion with plausible contact. We further demonstrate the advantage of our pipeline, by generating results for a various previously unexplored target objects, and also validating its advantage in merging multiple HOI motion concepts with known human motion from pre-trained MDM [92].
@@
In summary, our main contributions are as follows: (1) We present a generative motion model, DAViD, to learn Dynamic Affordance; (2) We present a pipeline to synthetically generate diverse 4D HOI samples leveraging video diffusion models; (3) We address generating various novel HOI motions by fusing multiple, separately trained LoRAs with known human motions from a pre-trained model. The results and our source code are publicly available.
@@ 2 Related Work
Reasoning Visual Affordance. Affordance is a concept defined by Gibson [23] as the set of possible interactions that an agent can perform within an environment. As the concept narrows to the primary agent (or humans) and the specific interaction target (or objects), the area of learning affordance mainly focuses on HOI. Specifically, followed by the intuition that the visual geometry of an object is closely related to its functionality, many studies leverage visual cues to reason about object’s affordance. The studies are subdivided into various areas, including affordance classification [97, 89, 83, 96, 60, 76, 43], affordance detection [1, 66, 64, 67, 82, 68, 17, 12], and affordance segmentation [79, 38, 58, 13]. Such traditional studies infer affordance for the given context (mostly provided as images), as a form of action labels [6, 32, 50], and contact (or interaction) regions in 2D [52, 2] or 3D [94, 114]. Later studies model static HOI patterns (e.g., contact) during interactions [15, 28, 39], going beyond the understanding of individual HOI situations. Recent studies synthesize human motion including HOI [111, 74, 104, 117, 54, 112, 108], showing potential to learn dynamic HOI patterns, but they focus on a text-driven approach, generating consistent human motions regardless of object categories (e.g., dragging a chair, holding a backpack). In contrast, our method focuses on learning rich, dynamic object-specific HOI patterns that are observable in real-world HOIs.
Learning from Synthesized Data. Generative models such as GANs [24] and diffusion models are used as a strategy to address the problem of insufficient real-world data in many fields. Starting from augmentation [8, 3, 93, 31, 107, 69], synthetic datasets are widely used in vision area including segmentation [61, 95, 123], classification [9, 59, 91], shape reconstruction [71], and neural rendering [29, 122]. While data synthesis is a powerful technique with broad applications, a key consideration when learning from generated data is its reliability compared to real-world data. Recently, a few studies in the HOI field [28, 39] leverage pre-trained 2D diffusion models to generate 2D HOI images, demonstrating the reliability of synthetic data to learn affordance knowledge. While following the approaches, we focus on learning dynamic patterns during HOI rather than learning static patterns.
@@ 3 Method
Our approach aims to learn the Dynamic Affordance of a target object from synthetically generated 2D videos, which are produced by a video diffusion model [44]. In the first stage, we generate diverse 4D HOI samples interacting with a given 3D object mesh by leveraging pre-trained video models (Sec. 3.1). Then, our method models the collected 4D HOI samples into our DAViD, which consists of (1) a motion diffusion model along with LoRA, and (2) an object pose diffusion model. The first module synthesizes feasible human motions for the target object, and the second module synthesizes object motions paired with previously generated human motion, using our novel sampling technique, Temporal Guidance Sampling (TGS) (Sec 3.2). An overview of our method is shown in Fig. 3.
@@
2D HOI Video Synthesis. Given an input 3D object mesh, we produce multiple 2D HOI videos capturing human-object interactions with the object. We first render the object from cameras with multiple viewpoints. The intrinsic parameter is fixed to match the same perspective camera model and parameters of the world-grounded human mesh recovery method (GVHMR [85]) used in our later 3D lifting process, ensuring a consistent camera setup. For the extrinsic parameters, we consider diverse viewpoints based on the object’s type, either ground-placed or portable objects. For relatively stationary objects (e.g., cart), we consider horizontal viewpoint variations centered on the object, while for the portable objects (e.g., trumpet) we sample viewpoints within a specified range from all view angles. The object type is automatically inferred via a vision-language model [70], by querying with the rendered image with a prompt: “Is the object plausibly on the ground while being used?”.
Once the object is rendered from each view, we synthesize a realistic human naturally interacting with it using an off-the-shelf pre-trained 2D image diffusion model [47]. Unlike the previous similar approaches [45, 116, 39] that use inpainting by providing a mask region to a model, we use ControlNet [120] to synthesize the entire image, including the human and the plausible background, by using edge cues of rendered object [7] as the condition. Our primary objective is to synthesize both plausible human-object interaction for the given object viewpoint and a realistic background, which is important for estimating dynamic cues and camera poses in a global coordinate. After synthesizing HOI images from several views, we refine the results by filtering out low-quality or undesirable images. The filtering process can be performed efficiently with minimal human effort, requiring only about three minutes per category.
@@
Figure 3: Overview. Our method consists of two parts: (1) Generating 4D HOI Samples and (2) Learning Dynamic Affordance from Generated 4D HOI Samples. First, we create 2D HOI videos, and generate 4D HOI samples with our uplifting pipeline. Then, the generated 4D HOI samples are used to train DAViD, learning the patterns of human motion and object pose.
Uplifting for 4D HOI Sample Generations. For each 2D HOI video $\mathcal{V}$ 222we drop the video index $d$ for simplicity., we uplift 3D human motions and 3D object movements to generate 4D HOI samples. Our key strategy is to uplift each component (human, object) by finding each motion relative to each camera, and co-locate them in the common camera coordinate shared by both components.
For the object side, we use the Perspective-n-Points (PnP) [21, 51] algorithm to find the relative camera pose with respect to the given 3D object mesh. Specifically, we aim to compute the extrinsic camera poses $\{\bm{R}_{o}^{f},\bm{t}_{o}^{f}\}_{f=1}^{F}$, where the one for the first frame $f=1$ is already known. We first identify visible object mesh vertices $\{\mathbf{V}_{o}\}_{\text{vis}}$ via raycasting [78] for the first frame $f=1$, where $\mathbf{V}_{o}\in\mathbb{R}^{3}$ is an object vertex. The 2D projections of the visible vertices at the first frame is denoted as $\{\bm{v}_{o}^{f=1}\}_{\text{vis}}$, where $\bm{v}_{o}^{f=1}\in\mathbb{R}^{2}$. By applying an off-the-shelf 2D video tracker [37, 36], we obtain the corresponding 2D points $\bm{v}_{o}^{f}$ in the remaining frames $f>1$. As we have the correspondences between 3D object vertices and matched 2D points for every frame, we apply PnP [21, 51] to compute the camera poses at frame $f$ with respect to the object mesh in its canonical space:
@@
For the human side, we leverage world-grounded human motion recovery models, denoted as $\mathbf{F}_{\text{human}}$ to uplift human motions and corresponding camera motions into 3D from the synthesized 2D HOI videos. We use GVHMR [85], except in the cases where human translation occurs without walking motions (e.g., riding a motorcycle). In such cases, we use TRAM [103], as GVHMR [85] typically fails to predict plausible camera motion under these conditions. Therefore,
@@
where $\bm{\theta}^{f}\in\mathbb{R}^{21\times 3}$, $\bm{\beta}^{f}\in\mathbb{R}^{10}$, $\bm{\phi}^{f}\in\text{SO(3)}$, $\bm{\tau}^{f}\in\mathbb{R}^{3}$ are the predicted SMPL-X [72] parameters of human pose, shape, root joint’s rotation, and translation for $f$-th frame, respectively. The $\bm{R}_{h}^{f}\in\text{SO(3)}$, $\bm{t}_{h}^{f}\in\mathbb{R}^{3}$ are the extrinsic parameters of estimated camera poses. Note that all of the parameters are defined in a world coordinate for the human mesh space, which is different from object mesh space.
@@
where $\bm{V}_{h}^{f}\in\mathbb{R}^{3}$ is a human 3D vertex in its global coordinate after applying all SMPL parameters, and $\bm{\hat{V}}_{o}^{f}$ and $\bm{\hat{V}}_{h}^{f}$ are the human and object 3D vertices in the shared camera coordinate. To this end, we align the object-side cameras to the human-side cameras, bringing cameras, humans, and objects into a unified coordinate system of human side.
@@
[…]
@@
where $\bm{\hat{V}}_{h,i}$ and $\bm{\hat{V}}_{o,i}$ are the $i$-th human and object vertices in the camera coordinate at the first frame respectively, and $\bm{P}_{h,\text{3D}}$, $\bm{P}_{o,\text{3D}}$ are the corresponding 3D points by back-projecting 2D metric depths (computed by metric depth model [5]) into the camera coordinate. Note that the indices $i,j$ are from the set of visible vertex indices of the human and object at the first frame.
@@ 3.2 Learning Dynamic Affordance
Human Motion Synthesis with LoRA for MDM. To capture human motion patterns interacting with a specific object, we insert a LoRA [33] layer into a pre-trained MDM [92] model and train it for each object category. By leveraging LoRA [33], our model can learn previously unseen human motion patterns with a limited amount of training samples, keeping the original motion synthesis quality of the pre-trained model. Interestingly, we find that merging LoRA [33] layers trained across different object categories enables our model to generate human motion interacting with multiple objects in a zero-shot manner (see Fig. 5 (c)).
@@ 4.2 Baselines and Metric
Even though our goal is to learn dynamic HOI patterns for various objects, which is slightly different from typical text-driven HOI motion synthesis [16, 74, 112, 117, 111, 106], we conduct quantitative comparisons on existing datasets for evaluation. Specifically, we compare DAViD against OMOMO [53] and CHOIS [54] on FullBodyManipulation Dataset [53], which provide available code and evaluate the quality of HOI, unlike other stuides [16, 74, 106] that focus on evaluating human quality solely. However, OMOMO [53] and CHOIS [54] generate human motion conditioned on object motion (or waypoints), making a direct comparison with ours challenging. To address this, we compare the HOI motion generated by baselines using object motion (or waypoints) condition with the HOI motion generated by DAViD using human motion condition. Each condition is extracted from the test dataset of FullBodyManipulation [53], enabling a comparison with the ground truth. Due to the different conditions, we compare only the quality of HOI for the fairness, specifically the precision, recall, and F1 score of the contact, following baselines [53, 54]. As DAViD performs modeling of object poses during interaction, we evaluate by cropping each test dataset sequence from the start of the interaction (contact start) to the end (contact end). Additionally, to assess the human motion quality, we report the Foot Sliding (FS) and foot height (in centimeters) of the human motion generated from DAViD and the baselines, both trained on FullBodyManipulation Dataset [53].